OpenStudy (anonymous):

What is the length of the line segment with endpoints (-8,6) and (-10,-1)? A. 53 B. sq root 373 C. sq root 53 D. 373

OpenStudy (anonymous):

tell me the 1st step

OpenStudy (anonymous):

The distance formula gets you the answer. The difference in x is 2 and the difference in y is 7. So, the distance between the points is: \[ \sqrt {2^2 + 7^2} = \sqrt {4 + 49} = \boxed {\sqrt{53}}. \]A

OpenStudy (anonymous):

thx. sorry for not replying. i had to feed my baby. lol

OpenStudy (anonymous):

lol how old r u

OpenStudy (anonymous):

16